OpenStudy (b77w):

Simplify: ((12x^-3)/(y^4))*(((y^-2x^2)^-1)/(15x^-2))

OpenStudy (unklerhaukus):

\[\frac{12x^{-3}}{y^4}\times \frac{(y^{-2}x^2)^{-1}}{15x^{-2}}\]

OpenStudy (unklerhaukus):

for the term in the brackets that its to the power of negative one, \[(y^{-2}x^2)^{-1}\] distribute the index, like this \[(a^nb^n)^p=a^{n\times p}b^{m\times p}\]

OpenStudy (unklerhaukus):

can you do that?

OpenStudy (b77w):

y^2x^-2

OpenStudy (b77w):

I got 4/(3xy^2)

OpenStudy (unklerhaukus):

well you got this bit right \((y^{−2}x^2)^{−1}=y^2x^{-2}\)\(\checkmark\)

OpenStudy (unklerhaukus):

your final result is close but not quite right

OpenStudy (unklerhaukus):

...\[\frac{12x^{-3}}{y^4}\times \frac{y^{2}x^{-2}}{15x^{-2}}\]...